OpenStudy (word2):

Jim is designing a seesaw for a children's park. The seesaw should make an angle of 30° with the ground, and the maximum height to which it should rise is 1.5 meters What is the maximum length of the seesaw? 3.0 meters 3.5 meters 4.0 meters 4.5 meters

OpenStudy (word2):



OpenStudy (word2):

i think its 3.5 , am i correct?

OpenStudy (word2):

@Qwertty123. @SnuggieLad

OpenStudy (qwertty123.):

@Qwertty123

OpenStudy (roxy.girl):

I am not sure it it is..3.5 let me work this out.

OpenStudy (word2):

okay

OpenStudy (roxy.girl):

1.5/sin 30=c/sin 90 . 1.5m/0.5=c/1 . 3m=c . additional info: in 30° 60° 90° special triangle you can just apply the side ratio to lessen calculations: . if the given is short leg like in your problem: . short leg= x long leg =x√3 hypotenuse = 2x . if hypotenuse is given just divide it by 2 to find x then apply x to the legs... ex: hypo= 4 4=2x 4/2=x 2=x so short leg=2, long leg=2(√3) . check by pythagorean: 2^2 plus [2(√3)]^2=c^2 √(4 plus 12)=c √16=c 4=c . . if long leg is given just multiply it by (√3)/3 to find x ex: long leg= 9 9*(√3)/3= x 3(√3)=x . so short leg= 3(√3), hypo=6(√3) if you check long leg: 3(√3)*(√3)=9 ~ https://answers.yahoo.com/question/index?qid=20130101154201AAZ9Ob4

OpenStudy (word2):

that's kinda confusing , i get the answer but i still don't understand the problem